A coil 0f 200 turns is wound uniformly over wooden ring having a mean circum ference of 600mm and a uniformly cross the sectional area of 500 mm^2 IF the current through the coil is 4A calculate
1)the magnetic field strength
2)the flux density
3)the total flux
1
Expert's answer
2011-06-10T05:32:24-0400
To find the flux of field there is a formula
B=μ *N *I/(2*π*r)

r- the mean radius. if the mean circumference s=0.6 m, the the mean radius
r=s/(2*π)=0.1
the value of munot is constant: μ=4(3.142)x10-7 web/amp m
Hence we have
B=4*3.142x10-7 * 200*4/(2*3.142*0.1)
B=1.6x10-3 weber/m2 or tesla
So we get magnetic field
For finding the magnetic flux density there is a formula

magnetic flux=magnetic field * Area

magnetic flux=1.6x10-3 * 0.5x10-3
magnetic flux=0.8x10-6 webers
Now flux density
flux density=magnetic flux/area
flux density=1.6x10-3 web/m2